# Proportionality for Constrained Public Decisions

### v1

@@ 3 Proportionality via Justified Representation
Suppose there are two issues =$\mathcal{I}=\{a_{1},a_{2}\}$ with constraint $\mathcal{C}=\{(1,0),(0,1)\}$. Then suppose there are two voters $N=\{1,2\}$ with ballots $\boldsymbol{b}_{1}=(1,0)$, and $\boldsymbol{b}_{2}=(0,1)$ (note that voters $1$ and $2$ are both, on their own, cohesive groups). Here, both aforementioned EJR interpretations require each voter to obtain at least $1$ in satisfaction, i.e., deciding half of the two issues at hand. However, there exists no feasible outcome that provides agreement-EJR or cohesiveness-EJR as one voter $i\in\{1,2\}$ will have satisfaction $u_{i}(\boldsymbol{w})=0$ for any outcome $\boldsymbol{w}\in\mathcal{C}$.
@@
Example 1 makes clear an issue that we must take into account when defining proportionality properties when there are constraints. That is, a voter group that is an $\alpha$ fraction of the population may lay claim to deciding an $\alpha$ fraction of the issues, but in doing so, they may be resolving, or influencing the decision on, a larger portion of the issues than they are entitled to.
In doing so, we look for meaningful ways to identify, given an outcome $\boldsymbol{w}$, those voter groups that are underrepresented and can justifiably complain at the selection of outcome $\boldsymbol{w}$.
The latter is formalised by the following definition which we use to identify the voter group whose displeasure is justified. Specifically, these are groups that can propose an alternative, feasible outcome $\boldsymbol{w}^{*}$ that yields greater satisfaction for each group member.
Definition 1 ($(S,\boldsymbol{w})$-deviation).
@@
Given election instance $(\boldsymbol{B},\mathcal{C})$ and outcome $\boldsymbol{w}\in\mathcal{C}$, a set of voters $N^{\prime}\subseteq N$ has an $(S,\boldsymbol{w})$-deviation if $\emptyset\neq S\subseteq\mathcal{I}$ is a set of issues such that all of the following hold:
@@
  • $S\subseteq\text{\rm Agr}(\boldsymbol{b}_{i},\boldsymbol{b}_{j})\text{ for all }i,j\in N^{\prime}$ (the voters agree on the decisions on all issues in $S$).
@@
  • $S\subseteq\mathcal{I}\setminus{\text{\rm Agr}(\boldsymbol{b}_{i},\boldsymbol{w})}$ for all $i\in N^{\prime}$
    (the voters disagree with outcome $\boldsymbol{w}$’s decisions on all issues in $S$).
@@
Intuitively, given an outcome $\boldsymbol{w}$, a voter group having an $(S,\boldsymbol{w})$-deviation indicates the presence of another feasible outcome $\boldsymbol{w}^{*}\neq\boldsymbol{w}$ where every group member would be better off. Thus, our goal in providing a fair outcome reduces to finding an outcome where every group of voters that has an $(S,\boldsymbol{w})$-deviation is sufficiently represented. We shall use this $(S,\boldsymbol{w})$-deviation notion to convert proportionality axioms from unconstrained settings to axioms that deal with constraints. But first, we look at the following computational question associated with $(S,\boldsymbol{w})$-deviations: given an election instance $(\boldsymbol{B},\mathcal{C})$ and an outcome $\boldsymbol{w}\in\mathcal{C}$, the problem is to find all groups of voters with an $(S,\boldsymbol{w})$-deviation.
Proposition 1.
Given an election instance $(\boldsymbol{B},\mathcal{C})$ and an outcome $\boldsymbol{w}\in\mathcal{C}$, there exists an algorithm that finds all groups of voters $N^{\prime}$ such that there exists an $S\subseteq\mathcal{I}$ with $N^{\prime}$ having an $(S,\boldsymbol{w})$-deviation, that runs in $O(|\mathcal{C}|^{2}mn)$ time.
Proof.
Take $(\boldsymbol{B},\mathcal{C})$ and outcome $\boldsymbol{w}\in\mathcal{C}$. Consider the following algorithm that operates in $|\mathcal{C}|$ rounds, assessing an outcome $\boldsymbol{w}\in\mathcal{C}$ in each round (with each outcome assessed once throughout): at each round for an outcome $\boldsymbol{w}\in\mathcal{C}$, iterate through all other outcomes $\boldsymbol{w}^{*}\neq\boldsymbol{w}\in\mathcal{C}$; fix $S$ to be the issues that $\boldsymbol{w}$ and $\boldsymbol{w}^{*}$ disagree on; in at most $mn$ steps, it can be checked if there is a set of voters that agree with $\boldsymbol{w}^{*}$ on all issues in $S$ which verifies the existence of a voter group $N^{\prime}$ with an $(S,\boldsymbol{w})$-deviation; keep track of all such groups $N^{\prime}$; if all outcomes have been assessed, terminate, otherwise, move to the next outcome. This algorithm takes $O(|\mathcal{C}|^{2}mn)$ time to complete in the worst case, which is polynomial in the input size given our assumptions. ∎
We offer the following remark in regards to the nature of Proposition 1.
Remark 2.
Proposition 1 can be seen as positive whenever the constraint $\mathcal{C}$ under consideration is ‘not too large’. Such an assumption is reasonable for many real-life examples. Consider the quite general, collective task of selecting the features of some product. Our running example of the logo design is an instance of this. Other applicable scenarios include choosing the technical features of a shared computer or the items to be placed in an organisation’s common area. In many cases, factors such as a limited budget (or limited space in the case of the common area) may result in very few feature combinations being feasible for said product. These are natural scenarios where we may encounter a ‘small’ constraint (according to our definition) with respect to the number of issues at hand and the size of their domains.
Our goal is to answer the following question: how much representation can we guarantee from some outcome $\boldsymbol{w}$, to a group of voters that has an $(S,\boldsymbol{w})$-deviation and that qualifies as underrepresented?
@@ 4 Justified Representation with Cohesiveness
We now propose the following adaptations of cohesiveness-EJR to public decisions with constraints. To adapt cohesiveness-EJR, we adapt cohesiveness from multiwinner voting in a similar manner as done by Freeman et al. (2020). We say that a voter group is $T$-agreeing for some set of issues $T\subseteq\mathcal{I}$ if $T\subseteq\text{\rm Agr}(\boldsymbol{b}_{i},\boldsymbol{b}_{j})$ holds for all voters $i,j\in N^{\prime}$ and then we define cohesiveness as the following:
@@
Using $T$-cohesiveness, we can define EJR for public decisions with constraints (Freeman et al., 2020).
@@
Given an election $(\boldsymbol{B},\mathcal{C})$, an outcome $\boldsymbol{w}$ provides $\text{cohEJR}_{\mathcal{C}}$ if for every $T$-cohesive group of voters $N^{\prime}\subseteq N$ for some $T\subseteq\mathcal{I}$ with an $(S,\boldsymbol{w})$-deviation for some $S\subseteq T$, there exists a voter $i\in N^{\prime}$ such that $u_{i}(\boldsymbol{w})\geqslant|T|$.
Intuitively, $\text{cohEJR}_{\mathcal{C}}$ deems an outcome to be unfair if there exists a $T$-cohesive voter group with $(i)$ none of its group members having at least $|T|$ in satisfaction, and $(ii)$ ‘flipping’ outcome $\boldsymbol{w}$’s decisions on some of the issues in $T$ leads to some other feasible outcome.
We have the following result that can be interpreted as positive when the size of $\mathcal{C}$ is ‘not too large’.
@@
Given an election instance $(\boldsymbol{B},\mathcal{C})$ and an outcome $\boldsymbol{w}\in\mathcal{C}$, there exists an algorithm that decides in $O((\max_{t\in[m]}|D_{t}|)^{m}|\mathcal{C}|^{3}mn)$ time whether outcome $\boldsymbol{w}$ provides $\text{cohEJR}_{\mathcal{C}}$.
@@
From Proposition 1 we know that, given an outcome $\boldsymbol{w}$, we can find all groups with some $(S,\boldsymbol{w})$-deviation for some $S\subseteq\mathcal{I}$ in $O(|\mathcal{C}|^{2}mn)$ time. There can be at most $(\max_{t\in[m]}|D_{t}|)^{m}(|\mathcal{C}|-1)$ such groups (recall that $\max_{t\in[m]}|D_{t}|$ is the maximal size of any issue’s domain). Then, for each group $N^{\prime}$ with an $(S,\boldsymbol{w})$-deviation, we can check their size in polynomial time and thus verify whether they are a $T$-cohesive with $S\subseteq T$, and if so, we can check if there exists any voter $i\in N^{\prime}$ with $u_{i}(\boldsymbol{w})>|T|$. ∎
Now, Chandak et al. (2024) have already shown that, in general, cohesiveness-EJR is not always satisfiable in their sequential decisions model. This negative result carries over to the unconstrained public-decision setting. Although we shall, in the sections to follow, analyse the extent to which we can achieve positive results with cohesiveness-EJR in our constrained setting, this negative result motivates the study of the following weaker axiom—which is an adaptation of the multiwinner JR axiom—that can always be satisfied in the public-decision setting without constraints (Bulteau et al., 2021; Chandak et al., 2024; Freeman et al., 2020).
@@
Given an election instance $(\boldsymbol{B},\mathcal{C})$, an outcome $\boldsymbol{w}$ provides $\text{cohJR}_{\mathcal{C}}$ if for every $T$-cohesive group of voters $N^{\prime}\subseteq N$ for some $T\subseteq\mathcal{I}$ with an $(S,\boldsymbol{w})$-deviation for some $S\subseteq T$ where $|S|=|T|=1$, there exists a voter $i\in N^{\prime}$ such that $u_{i}(\boldsymbol{w})\geqslant 1$.
Unfortunately, when considering arbitrary constraints, even $\text{cohJR}_{\mathcal{C}}$ cannot always be achieved. Note that this even holds for binary election instances.
@@
Consider the binary election instance with issues =$\mathcal{I}=\{a_{1},a_{2}\}$ and a constraint $\mathcal{C}=\{(0,1),(0,0)\}$. Suppose that $N=\{1,2\}$, where $\boldsymbol{b}_{1}=(1,1)$ and $\boldsymbol{b}_{2}=(1,0)$. Note that for both outcomes $\boldsymbol{w}\in\mathcal{C}$, one voter will have satisfaction of 00 while being a $T$-cohesive group with an $(S,\boldsymbol{w})$-deviation for $|S|=|T|=1$. As each voter is half of the population, they may ‘flip’ issue $a_{2}$ to deviate towards the alternative feasible outcome, which provides them greater satisfaction than the current one. ∎
Let us now restrict the constraints that we consider. To do so, we introduce notation for the fixed decisions for a set of outcomes $C\subseteq\mathcal{C}$, which are the issues in $\mathcal{I}$ whose decisions are equivalent across all the outcomes in $C$. For a set of outcomes $C\subseteq\mathcal{C}$, we represent this as:
$\mathcal{I}_{\text{fix}}(C)=\{a_{t}\in\mathcal{I}\mid\text{there exists some }d\in D_{t}\text{ such that }w_{t}=d\text{ for all }\boldsymbol{w}\in C\}.$
@@
We say a constraint $\mathcal{C}$ has the NFD property if $\mathcal{I}_{\text{fix}}(\mathcal{C})=\emptyset$ holds for $\mathcal{C}$.
Remark 3.
At first glance, this NFD property seems more than a reasonable requirement but rather a property that should be assumed to hold by default. We argue however, that by doing so, we will neglect election instances where decisions that are fixed from the get-go may contribute to the satisfaction of voters and, specifically for our goal, these fixed decisions may aid in giving the voters their fair, proportional representation. It is this reason, why we did not restrict ourselves to election instances where the NFD property holds.
Now, we show that with the NFD property, the $\text{cohEJR}_{\mathcal{C}}$ axiom can always be provided, albeit only for ‘small’ election instances. We begin with cases where the number of feasible outcomes is limited to two.
@@
Take some feasible outcome $\boldsymbol{w}\in\mathcal{C}$. Observe that when $|\mathcal{C}|=2$, if property NFD holds, then the two feasible outcomes differ on the decisions of all issues. Thus, it is only possible for $T$-cohesive groups with an $(S,\boldsymbol{w})$-deviation for $|S|\leqslant|T|=m$ to have an allowable deviation from $\boldsymbol{w}$ to the only other feasible outcome. This means only the entire voter population have the potential to deviate. And if such deviation to $\boldsymbol{w}^{\prime}$ exists, then outcome $\boldsymbol{w}^{\prime}$ sufficiently represents the entire voter population. ∎
@@
The case for $m=1$ is trivially satisfied so we present the proof as two separate cases where the number of issues is either $m=2$ or $m=3$.
Case $\boldsymbol{m=2}$: Observe that for two issues (i.e., $m=2$) there are $7$ possible constraints $\mathcal{C}$ satisfying the NFD property. Take one such $\mathcal{C}$ and a feasible outcome $\boldsymbol{w}=(d_{x},d_{y})\in\mathcal{C}$ where $d_{x},d_{y}\in\{0,1\}$. Let us consider now groups of voters with an $(S,\boldsymbol{w})$-deviation over some set of issues $S\subseteq T$ who are witness to a violation of $\text{cohEJR}_{\mathcal{C}}$. As $m=2$, the agreement among voters and the deviation may concern at most two issues, i.e., $|S|,|T|\in\{1,2\}$.

### v2

@@ 3.1 Proportionality via Justified Representation
Suppose there are two binary issues $\mathcal{I}=\{a_{1},a_{2}\}$ with constraint $\mathcal{C}=\{(1,0),(0,1)\}$. Then suppose there are two voters $N=\{1,2\}$ with ballots $\boldsymbol{b}_{1}=(1,0)$, and $\boldsymbol{b}_{2}=(0,1)$. Here, both aforementioned EJR interpretations require each voter to obtain at least $1$ in satisfaction, i.e., deciding half of the two issues at hand. This would be possible by selecting outcomes $(1,1)$ or $(0,0)$, which are however not feasible according to the constraints. Thus, there exists no feasible outcome that provides agreement-EJR or cohesiveness-EJR as one voter $i\in\{1,2\}$ will have satisfaction $u_{i}(\boldsymbol{w})=0$ for any outcome $\boldsymbol{w}\in\mathcal{C}$.
@@
Example 2 shows that a voter group that is an $\nicefrac{{1}}{{2}}$-fraction of the population may lay claim to deciding an $\nicefrac{{1}}{{2}}$-fraction of the issues, but in doing so, they may be resolving, or influencing the decision on, a larger portion of the issues than they are entitled to. This is an issue we must take into account when defining proportionality when there are constraints.
3.2 Constrained Deviations
To define proportionality axioms that accommodate constraints, we seek meaningful ways to identify, given an outcome $\boldsymbol{w}$, those voter groups that are happier with the selection of an outcome that is different from $\boldsymbol{w}$. We formalize this by the notion of deviation, in line with related work (Aziz et al., 2017; Masarík et al., 2024; Haret et al., 2024).
Definition 1 ($(S,\boldsymbol{w},\mathcal{C})$-deviation).
@@ 3.2 Constrained Deviations
Given election instance $(\boldsymbol{B},\mathcal{C})$ and outcome $\boldsymbol{w}\in\mathcal{C}$, a set of voters $N^{\prime}\subseteq N$ has an $(S,\boldsymbol{w},\mathcal{C})$-deviation if $\emptyset\neq S\subseteq\mathcal{I}$ is a set of issues such that all of the following hold:
@@
  • $S\subseteq\text{\rm Agr}(\boldsymbol{b}_{i},\boldsymbol{b}_{j})\text{ for all }i,j\in N^{\prime}$ (the voters agree with each other on the decisions on all issues in $S$).
@@
  • $S\subseteq\mathcal{I}\setminus{\text{\rm Agr}(\boldsymbol{b}_{i},\boldsymbol{w})}$ for all $i\in N^{\prime}$ (each voter disagrees with outcome $\boldsymbol{w}$’s decisions on all issues in $S$).
@@
We provide an example of a deviation as defined above with the following example.
Example 3.
Suppose there are three binary issues $\mathcal{I}=\{a_{1},a_{2},a_{3}\}$ with constraint $\mathcal{C}=\{(0,1,1),(1,0,1)\}$, and three voters $N=\{1,2,3\}$ with ballots $\boldsymbol{b}_{1}=(1,0,0)$, $\boldsymbol{b}_{2}=(1,0,0)$ and $\boldsymbol{b}_{3}=(0,0,0)$. Take outcome $\boldsymbol{w}=(0,1,1)$. Voters $N^{\prime}=\{1,2\}$ agree on issues in $S=\{a_{1},a_{2}\}$ and they disagree with $\boldsymbol{w}$ on the decisions for issues in $S$ and ${\boldsymbol{w}}[{S\leftarrow\boldsymbol{b}_{1}}]={\boldsymbol{w}}[{S\leftarrow\boldsymbol{b}_{2}}]=(1,0,1)\in\mathcal{C}$. Hence, voters $1$ and $2$ have an $(S,\boldsymbol{w},\mathcal{C})$-deviation to outcome $(1,0,1)$.
Intuitively, given an outcome $\boldsymbol{w}$, a voter group having an $(S,\boldsymbol{w},\mathcal{C})$-deviation indicates the presence of another feasible outcome $\boldsymbol{w}^{*}\neq\boldsymbol{w}$ where every group member would be better off. Thus, our goal in providing a fair outcome reduces to finding an outcome where every group of voters that has an $(S,\boldsymbol{w},\mathcal{C})$-deviation is sufficiently represented.
@@ 4 Justified Representation with Cohesiveness
To adapt cohesiveness-EJR to public decisions with constraints, we start from the notion of cohesiveness introduced in proportionality studies in the committee elections literature. We say that a voter group $N^{\prime}$ is $T$-agreeing for some set of issues $T\subseteq\mathcal{I}$ if $T\subseteq\text{\rm Agr}(\boldsymbol{b}_{i},\boldsymbol{b}_{j})$ holds for all voters $i,j\in N^{\prime}$ and then we define cohesiveness as the following:
@@
Using $T$-cohesiveness, we can define our version of EJR for public decisions with constraints.
@@
Given an election $(\boldsymbol{B},\mathcal{C})$, an outcome $\boldsymbol{w}$ provides $\text{cohEJR}_{\mathcal{C}}$ if for every $T$-cohesive group of voters $N^{\prime}\subseteq N$ for some $T\subseteq\mathcal{I}$ with an $(S,\boldsymbol{w},\mathcal{C})$-deviation for some non-empty $S\subseteq T$, there exists a voter $i\in N^{\prime}$ such that:
$u_{i}(\boldsymbol{w})\geqslant|T|.$
Intuitively, $\text{cohEJR}_{\mathcal{C}}$ deems an outcome to be unfair if there exists a $T$-cohesive voter group with $(i)$ none of its group members having at least $|T|$ in satisfaction, and $(ii)$ changing outcome $\boldsymbol{w}$’s decisions on some of the issues in $T$, specifically those in a subset $S\subseteq T$, leads to some other feasible outcome.
Example 4.
Suppose there are three issues $\mathcal{I}=\{a_{1},a_{2},a_{3}\}$ with domains $D_{t}=\{d_{1},d_{2},d_{3}\}$ for all $t\in\{1,2,3\}$. Then take the constraint to be $\mathcal{C}=\{(d_{1},d_{2},d_{3}),(d_{2},d_{1},d_{3})\}$. Then, suppose there are three voters $N=\{1,2,3\}$ with ballots $\boldsymbol{b}_{1}=(d_{1},d_{2},d_{1})$, $\boldsymbol{b}_{2}=(d_{1},d_{2},d_{2})$ and $\boldsymbol{b}_{3}=(d_{3},d_{3},d_{2})$. Note that voters $1$ and $2$ form a $T$-cohesive group for the set of issues $T=\{a_{1},a_{2}\}$ while voter group $\{2,3\}$ is a $T$-cohesive group for $T=\{a_{3}\}$. Take outcome $\boldsymbol{w}=(d_{2},d_{1},d_{3})$. Each of voters $N^{\prime}=\{1,2\}$ has a satisfaction of 0 and the group has an $(S,\boldsymbol{w},\mathcal{C})$-deviation (with $S=T$) to the alternative outcome $\boldsymbol{w}^{*}=(d_{1},d_{2},d_{3})$ thus, outcome $\boldsymbol{w}$ does not provide $\text{cohEJR}_{\mathcal{C}}$. For outcome $\boldsymbol{w}^{*}$, see that while the voters are a $T$-cohesive group for $T=\{a_{3}\}$, they do not have an $(S,\boldsymbol{w}^{*},\mathcal{C})$-deviation to a feasible outcome. Thus, it follows that outcome $\boldsymbol{w}^{*}$ provides $\text{cohEJR}_{\mathcal{C}}$.
We now discuss our proposal in the context of recent adaptations of proportionality axioms to constraints. We start with the work of Masarík et al. (2024), precisely with the axioms of Base EJR and EJR (Definitions 2 and 4 in the cited paper). Consider the election instance in Example 4. In this example, we argue that voters $\{1,2\}$ is $T$-cohesive for issues $T=\{a_{1},a_{2}\}$ and is witness to a violation of $\text{cohEJR}_{\mathcal{C}}$ by the outcome $\boldsymbol{w}=(d_{2},d_{1},d_{3})$ (at least one of them deserves $|T|=2$ in satisfaction). However, in the context of the Base EJR axiom of Masarík et al. (2024), voter group $\{1,2\}$ would not qualify as a group deserving of representation. To see this, we find a partial outcome within the feasible ones in $\mathcal{C}$, namely $(d_{2},\_,d_{3})$, such that both of the following hold: $(1)$ the group cannot ‘complete’ the partial outcome with decisions that they agree on, and $(2)$ the group size is not strictly larger than $n\cdot\nicefrac{{|T|}}{{(2+|T|)}}$ (where $2$ is the number of decided issues in the partial outcome $(d_{2},\_,d_{3})$). The same holds for the stronger EJR axiom by Masarík et al. (2024), which asks to consider only partial outcomes agreeing with $\boldsymbol{w}=(d_{2},d_{1},d_{3})$ ($(d_{2},\_,d_{3})$ still qualifies). These considerations illustrate that the axioms of Masarík et al. (2024) are more demanding of voter groups when assessing whether they deserve representation. We argue that our proposal of $\text{cohEJR}_{\mathcal{C}}$ axiom (and other notions to follow) are more intuitive and appropriately lenient when identifying groups worthy of representation but we do note that this makes our notion much more challenging to satisfy. Similar reasoning applies to Restrained EJR Mavrov et al. (2023). We direct readers to (Masarík et al., 2024) for an in-depth discussion (Section $3.2$ of their paper) on the similarities between Restrained EJR and Base EJR.
Next, we ask how computationally demanding it is to check whether some outcome provides $\text{cohEJR}_{\mathcal{C}}$. But first, we look at the following computational question associated with $(S,\boldsymbol{w},\mathcal{C})$-deviations: given an election instance $(\boldsymbol{B},\mathcal{C})$ and an outcome $\boldsymbol{w}\in\mathcal{C}$, the problem is to find all groups of voters with an $(S,\boldsymbol{w},\mathcal{C})$-deviation.
@@
Given an election instance $(\boldsymbol{B},\mathcal{C})$ and an outcome $\boldsymbol{w}\in\mathcal{C}$, there exists an algorithm that finds all groups of voters $N^{\prime}$ such that there exists an $S\subseteq\mathcal{I}$ with $N^{\prime}$ having an $(S,\boldsymbol{w},\mathcal{C})$-deviation, that runs in $O((\max_{t\in[m]}|D_{t}|)^{m}|\mathcal{C}|^{2}mn)$ time.
@@
Take $(\boldsymbol{B},\mathcal{C})$ and outcome $\boldsymbol{w}\in\mathcal{C}$. Consider the following algorithm that operates in $|\mathcal{C}|$ rounds, assessing an outcome $\boldsymbol{w}\in\mathcal{C}$ in each round (with each outcome assessed once throughout): at each round for an outcome $\boldsymbol{w}\in\mathcal{C}$, iterate through all other outcomes $\boldsymbol{w}^{*}\neq\boldsymbol{w}\in\mathcal{C}$; for all sets of issues that $\boldsymbol{w}$ and $\boldsymbol{w}^{*}$ disagree on, we fix $S$ to be such a set (there are at most $(\max_{t\in[m]}|D_{t}|)^{m}$ of these sets where $\max_{t\in[m]}|D_{t}|$ is the maximal size of any issue’s domain); in at most $mn$ steps, it can be checked if there is a set of voters that agree with $\boldsymbol{w}^{*}$ on all issues in $S$ which verifies the existence of a voter group $N^{\prime}$ with an $(S,\boldsymbol{w},\mathcal{C})$-deviation; keep a count of all such groups $N^{\prime}$; if all outcomes have been assessed, terminate and return the number of groups identified that have a $(S,\boldsymbol{w},\mathcal{C})$-deviation, otherwise, move to the next outcome. This algorithm takes $O((\max_{t\in[m]}|D_{t}|)^{m}|\mathcal{C}|^{2}mn)$ time to complete in the worst case. ∎
Given this result, we continue with the following computational result regarding the checking of whether $\text{cohEJR}_{\mathcal{C}}$ is satisfied by an outcome of some election instance.
Proposition 2.
Given an election instance $(\boldsymbol{B},\mathcal{C})$ and an outcome $\boldsymbol{w}\in\mathcal{C}$, there exists an algorithm that decides in $O(2(\max_{t\in[m]}|D_{t}|)^{m}|\mathcal{C}|^{3}mn)$ time whether outcome $\boldsymbol{w}$ provides $\text{cohEJR}_{\mathcal{C}}$.
Proof.
From Proposition 1 we know that, given an outcome $\boldsymbol{w}$, we can find all groups with some $(S,\boldsymbol{w},\mathcal{C})$-deviation for some $S\subseteq\mathcal{I}$ in $O(\max_{t\in[m]}|D_{t}|)^{m}|\mathcal{C}|^{2}mn)$ time. There can be at most $(\max_{t\in[m]}|D_{t}|)^{m}(|\mathcal{C}|-1)$ such groups (recall that $\max_{t\in[m]}|D_{t}|$ is the maximal size of any issue’s domain). Then, for each group $N^{\prime}$ with an $(S,\boldsymbol{w},\mathcal{C})$-deviation, we can check their size in polynomial time and thus verify whether they are $T$-cohesive with $S\subseteq T$, and if so, we can check if there exists any voter $i\in N^{\prime}$ with $u_{i}(\boldsymbol{w})>|T|$. ∎
This result is similarly negative in comparison to Proposition 2 as it is exponential in the number of issues $m$. Moving on, we mention that Chandak et al. (2024) have already shown that, in general, cohesiveness-EJR is not always satisfiable in their sequential decisions model. This negative result carries over to the unconstrained public-decision setting and thus, to the constrained setting as well. Although we shall, in the sections to follow, analyse the extent to which we can achieve positive results with cohesiveness-EJR in our constrained setting, this negative result motivates the study of the following weaker axiom—which is an adaptation of the JR axiom from committee elections—that can always be satisfied in the public-decision setting without constraints (Bulteau et al., 2021; Freeman et al., 2020; Chandak et al., 2024).
@@
Given an election instance $(\boldsymbol{B},\mathcal{C})$, an outcome $\boldsymbol{w}$ provides $\text{cohJR}_{\mathcal{C}}$ if for every $T$-cohesive group of voters $N^{\prime}\subseteq N$ for some $T\subseteq\mathcal{I}$ with an $(S,\boldsymbol{w},\mathcal{C})$-deviation for some $S\subseteq T$ where $|S|=|T|=1$, there exists a voter $i\in N^{\prime}$ such that: $u_{i}(\boldsymbol{w})\geqslant 1$.
This is a relaxation of $\text{cohEJR}_{\mathcal{C}}$ where the only voter groups that need be guaranteed any amount of satisfaction are those voter groups that are cohesive on, and have a feasible deviation with, a single issue. $\text{cohJR}_{\mathcal{C}}$ aims to ensure that such voter groups have at least one voter that agrees with the outcome’s decision on at least one issue.
Example 5.
Take the same election instance from Example 4 with three issues $\mathcal{I}=\{a_{1},a_{2},a_{3}\}$ with domains $D_{t}=\{d_{1},d_{2},d_{3}\}$ for all $t\in\{1,2,3\}$. Then take the constraint to be $\mathcal{C}=\{(d_{1},d_{1},d_{2}),(d_{2},d_{2},d_{2}),(d_{2},d_{2},d_{3})\}$. Then, suppose there are three voters $N=\{1,2,3\}$ with ballots $\boldsymbol{b}_{1}=(d_{1},d_{1},d_{1})$, $\boldsymbol{b}_{2}=(d_{1},d_{1},d_{2})$ and $\boldsymbol{b}_{3}=(d_{3},d_{3},d_{3})$. Take outcome $\boldsymbol{w}=(d_{2},d_{2},d_{2})$. The voter group $\{1,2\}$ is a $T$-cohesive group for the set of issues $T=\{a_{1},a_{2}\}$, both are not satisfied by any decisions by outcome $\boldsymbol{w}$ and they have an $(S,\boldsymbol{w},\mathcal{C})$-deviation (with $S=T$) to the alternative outcome $(d_{1},d_{1},d_{2})$ but we have that $|S|=|T|=2>1$ and thus, $\text{cohJR}_{\mathcal{C}}$ does not capture their complaint. In fact, the voter group $\{1,2\}$ does not have a feasible deviation of only a single issue for any of the outcomes in $\mathcal{C}$. So, this pair of voters are not witness to a violation of $\text{cohJR}_{\mathcal{C}}$. On the other hand, voter $3$ represents a $T$-cohesive group for $T=\{a_{3}\}$ and they have an $(S,\boldsymbol{w},\mathcal{C})$-deviation of size $|S|=|T|=1$ to the outcome $(d_{2},d_{2},d_{3})$. Thus, outcome $\boldsymbol{w}=(d_{2},d_{2},d_{2})$ does not provide $\text{cohJR}_{\mathcal{C}}$ but the outcome $\boldsymbol{w}=(d_{2},d_{2},d_{3})$ does.
The above example demonstrates that $\text{cohJR}_{\mathcal{C}}$ is a fairly tame demand. Unfortunately, when considering arbitrary constraints, even $\text{cohJR}_{\mathcal{C}}$ cannot always be achieved. Note that this even holds for binary election instances.
@@
Consider the binary election instance with issues $\mathcal{I}=\{a_{1},a_{2}\}$ and a constraint $\mathcal{C}=\{(0,1),(0,0)\}$. Suppose that $N=\{1,2\}$, where $\boldsymbol{b}_{1}=(1,1)$ and $\boldsymbol{b}_{2}=(1,0)$. Note that for both outcomes $\boldsymbol{w}\in\mathcal{C}$, one voter will have satisfaction of 0 while being a $T$-cohesive group with an $(S,\boldsymbol{w},\mathcal{C})$-deviation for $|S|=|T|=1$. As each voter is half of the population, they would have a deviation towards the alternative feasible outcome obtained by ‘flipping’ issue $a_{2}$, which provides them greater satisfaction than the current one. ∎
This result still holds if we restrict voters’ ballots to be consistent with $\mathcal{C}$ (see Proposition 6’s counterexample).
Moving on, we propose a restriction to the class of constraints that we consider, given the negative results for the general case. To do so, we introduce notation for the fixed decisions for a set of outcomes $C\subseteq\mathcal{C}$, which are the issues in $\mathcal{I}$ whose decisions are equivalent across all the outcomes in $C$. For a set of outcomes $C\subseteq\mathcal{C}$, we represent this as:
[…]
@@
A constraint $\mathcal{C}$ has the NFD property if $\mathcal{I}_{\text{fix}}(\mathcal{C})=\emptyset$ holds for $\mathcal{C}$.
While the NFD property seems rather natural, we argue that there are cases in which considering election instances where decisions that are fixed from the get-go may contribute to the satisfaction of voters and, specifically for our goal, these fixed decisions may aid in giving the voters their fair, proportional representation.
Now, we show that with the NFD property, the $\text{cohEJR}_{\mathcal{C}}$ axiom can always be satisfied, albeit only for election instances where either $|\mathcal{C}|=2$ (i.e., there are only two feasible outcomes) or $m\leqslant 3$ (i.e., the number of issues is lower than three). We begin with the (essentially) trivial case where the number of feasible outcomes is limited to two.
@@
Take some feasible outcome $\boldsymbol{w}\in\mathcal{C}$. Observe that when $|\mathcal{C}|=2$, if property NFD holds, then the two feasible outcomes differ on the decisions of all issues. Thus, it is only possible for $T$-cohesive groups with an $(S,\boldsymbol{w},\mathcal{C})$-deviation for $|S|\leqslant|T|=m$ to have an allowable deviation from $\boldsymbol{w}$ to the only other feasible outcome. This means only the entire voter population has the potential to deviate. And if such deviation to $\boldsymbol{w}^{\prime}$ exists, then outcome $\boldsymbol{w}^{\prime}$ sufficiently represents the entire voter population. ∎
@@
The case for $m=1$ is trivially satisfied (since with only one issue, we get there only being two feasible outcomes and thus, any possible deviating group has to be of size $n$) so we present the proof as two separate cases where the number of issues is either $m=2$ or $m=3$.
Case $\boldsymbol{m=2}$: Observe that for two binary issues (i.e., $m=2$), there are $16$ non-empty constraints $\mathcal{C}$, but only $7$ of them satisfy the NFD property. Take one such $\mathcal{C}$ and a feasible outcome $\boldsymbol{w}=(d_{1},d_{2})\in\mathcal{C}$ where $d_{1},d_{2}\in\{0,1\}$.
We have to show that for every $T$-cohesive groups of voters $N^{\prime}\subseteq N$ for some $T\subseteq\mathcal{I}$ with an $(S,\boldsymbol{w},\mathcal{C})$-deviation for some non-empty $S\subseteq T$, there exists a voter $i\in N^{\prime}$ such that $u_{i}(\boldsymbol{w})\geqslant|T|$. Since $|\mathcal{I}|=m=2$, the agreement among voters (i.e., the set $T$) and their potential deviation (i.e., the set $S$) may concern at most two issues: namely, $|S|,|T|\in\{1,2\}$.